OpenStudy (anonymous):

A student builds an experiment model rocket with a variable thrust engine. The test flight last for 5 seconds and the rocket's altitude (in meters) is given as s(t)=1/3(t)^(3)-2t^(2)+3t+2 a) Find the velocity and acceleration of the rocket at t=4 b)When is the particle at rest? c)When is the acceleration 0? d)Find the altitude and the velocity at the time(s) when the acceleration is 0 e)Find the total distance traveled in the first 5 seconds

OpenStudy (anonymous):

velocity is the derivative acceleration is the second derivative

OpenStudy (anonymous):

b)When is the particle at rest? set the derivative equal to zero and solve

OpenStudy (anonymous):

I just want to make sure I did it right, for a I got 3m/s and 4m/s^2 but I'm not sure if it's correct

OpenStudy (anonymous):

i don't know the derivative is \[v(t)=t^2-4t+3\] and \[a(t)=2t-4\]

OpenStudy (anonymous):

is \(v(4)=3\)?

OpenStudy (anonymous):

yup that's what I got and I got a(4)=4

OpenStudy (anonymous):

I think I got a and b but how about c?

OpenStudy (anonymous):

Solve for 0=2t-4 Get t=2

OpenStudy (anonymous):

ok thank you :) I'm kind of confused on d. How do you set it up?

OpenStudy (anonymous):

Well the original formula is the position aka altitude. So plug t=2 into that. The first derivative is the velocity so plug t=2 into that. That should be what you are looking for

OpenStudy (anonymous):

ok I got it thank you :)